OpenStudy (anonymous):

Analyze what I did incorrectly when I tried to find the angle between the two lines x+2y=8; x-2y=2

OpenStudy (anonymous):

\[\frac{.5--.5}{1-.5*-.5} = \frac{4}{5}\]

OpenStudy (anonymous):

\[\tan^{-1} {4/5} =.6747\]

OpenStudy (anonymous):

The book says it is .9373

OpenStudy (anonymous):

.927295 that I get.

OpenStudy (anonymous):

\[2 tan^-1(.5)\]

OpenStudy (anonymous):

formula is \[\huge \frac{m1-m2}{1+m1.m2}\]

OpenStudy (anonymous):

Oh, I subtracted!

OpenStudy (anonymous):

Thank you, lol.

OpenStudy (anonymous):

welcome dear